OpenStudy (anonymous):

will give medal and fan best answer :-) Please help me with 3 questions

OpenStudy (anonymous):

1.) Change 28 degrees to radians. Please show all work and reduce for full credit. (10 pts) 2.) What are the sine, cosine, and tangent of radians? Show any work that is needed to solve this please. (30 pts) 3.) If sin Θ = 3/5 and tan Θ > 0, what is the value of cos Θ? Please show all work for full credit. (10 pts.)

OpenStudy (anonymous):

@kropot72 @amistre64 @whpalmer4 @robtobey

OpenStudy (anonymous):

@secret66

OpenStudy (anonymous):

@ganeshie8 @texaschic101

OpenStudy (anonymous):

@whpalmer4 @Warriorz13

OpenStudy (warriorz13):

honestly not to disappoint you but i m not good at this kind of stuff

OpenStudy (anonymous):

thank you for your time

OpenStudy (warriorz13):

NP

OpenStudy (whpalmer4):

To convert degrees to radians, remember that \(360^\circ = 2\pi \text{ radians}\) and set up the proportion. \[\frac{x^\circ}{360^\circ} = \frac{y\text{ radians}}{2\pi\text{ radians}}\] Plug in \(x\) or \(y\) depending on which one you know, and solve for the other.

OpenStudy (whpalmer4):

Can't help you with the second question, because you didn't care enough to make sure that everything showed up!